OpenStudy (anonymous):

m/3 + 1/3= 2/3 How do I solve?

OpenStudy (anonymous):

Well if you notice, they have the same denominator, so add the numerators. it will give you: (m+1)/3= 2/3 Now you can multiply both sides by 3 and you are left with m+1=2 and so you get m=1

OpenStudy (texaschic101):

m/3 + 1/3 = 2/3 (subtract 1/3 from both sides) m/3 = 2/3 - 1/3 m/3 = 1/3 (now multiply the equation by 3) m = 1